# arXiv:2602.22626v2[cond-mat.str-el] 02 Mar 2026

### v2

@@ Abstract
A key limitation of conventional approaches is their formulation in terms of a small number of collective order-parameter fields [10, 11]. Such descriptions implicitly assume that the infrared behavior can be captured by a finite set of soft bosonic modes with local-in-time damping. However, strongly correlated quantum matter is fundamentally governed by Liouvillian evolution of the density matrix, where the full spectrum of relaxation eigenvalues determines long-time dynamics [12, 13, 14]. In more general, criticality need not correspond to softening of a single excitation mode but may instead involve reorganization of the entire relaxation-rate spectrum.
Rather than reducing the dynamics to a few collective coordinates, one must consider the possibility that the Liouvillian spectrum develops an extended continuum of near-marginal decay modes. Such spectral accumulation naturally generates a broad distribution of relaxation time scales and invalidates the assumption of Markovian coarse-graining near criticality. The relevant organizing structure of the infrared theory is then not a small set of overdamped bosons, but the time–scale density of states (TDOS) of collective relaxation modes.
@@
We demonstrate that a TDOS scaling $\rho(\lambda)\sim\lambda^{\alpha}$ directly classifies universality classes of quantum critical dynamics. While the conventional Hertz–Millis kernel emerges only for $\alpha=1$, a finite TDOS at the origin ($\alpha=0$) produces a marginal nonanalytic retarded kernel proportional to $\mathrm{sgn}(\omega)$, corresponding in the time domain to a universal long–memory response $K(t)\sim 1/t$. This memory–dominated infrared dynamics represents a breakdown of Markovian overdamped criticality and defines a distinct universality class governed by the structure of the relaxation spectrum.
@@
the dissipative part of the collective kernel follows directly from the spectral representation in Eq. (16),
@@
Here $\omega$ denotes the collective relaxation frequency arising from the projected Liouvillian dynamics, rather than a microscopic quasiparticle excitation energy.
It is therefore natural to write the long–wavelength dynamical response in the scaling form
@@
in direct analogy with the conventional Hertz–Millis representation [3, 4]. The TDOS scaling immediately implies the universal identification $s=\alpha$.
Under the renormalization–group rescaling $\mathbf{q}\to b\,\mathbf{q}$ and $\omega\to b^{z}\omega$, the spatial gradient term scales as $q^{2}\to b^{2}q^{2}$, while the dissipative term scales as $|\omega|^{s}\to b^{zs}|\omega|^{s}$. Requiring both contributions to remain of comparable order at criticality ($r=0$) fixes the dynamical exponent to
$z=\frac{2}{s}.$ (20)
In this framework the dissipation is generically non–Ohmic, $|\omega|^{s}$ with $s\neq 1$, reflecting the organization of the underlying relaxation spectrum. Sub–Ohmic memory ($s<1$) implies $z>2$, corresponding to critical slowing down stronger than ordinary diffusive dynamics, whereas super–Ohmic behavior ($s>1$) yields $z<2$. Infrared universality classes are thus classified directly by the low–energy structure of the relaxation-rate density.
In particular, when the TDOS remains finite at the origin,
$\rho(\lambda)\xrightarrow{\lambda\to 0}\rho_{0},$ (21)
the dissipative kernel simplifies to
@@
while the reactive part acquires a logarithmic infrared singularity,
@@
Collective infrared dynamics is therefore governed by the marginal non-Markovian kernel
@@
corresponding in the time domain to a universal long-memory response $K(t)\sim 1/t$, derived explicitly in Appendix C.
This flat-TDOS fixed point represents a breakdown of the conventional overdamped critical dynamics underlying Hertz–Millis theory. Rather than arising from a fast electronic bath coupled to a single soft order-parameter mode, collective relaxation is controlled by the self-organized accumulation of an extensive continuum of slow decay modes.
@@
The emergence of a slow–mode reservoir reorganizes the infrared structure of the Cooper pairing channel. In conventional quantum critical superconductivity, collective fluctuations act as overdamped bosonic mediators whose Ohmic dynamics renders the pairing interaction marginal. Here we show that when collective dynamics is instead governed by a flat relaxation–rate density of states, long–time memory effects transform the Cooper channel into a relevant infrared instability.
III.1 A. Marginal pairing from Ohmic collective dynamics
Near metallic quantum critical points, fermions couple to overdamped collective modes with retarded propagator
@@
The linear frequency dependence reflects Landau damping by the electronic continuum and produces a pairing kernel scaling as
$D(\omega)\sim\frac{1}{|\omega|}.$ (26)
The infrared shell correction to the dimensionless Cooper–channel coupling $g$ therefore reads
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}=g^{2}\ln b.$ (27)
Introducing the RG scale $l=\ln b$ yields
$\frac{dg}{dl}=Cg^{2},$ (28)
with $C>0$.
@@
The pairing interaction is thus marginal and grows only logarithmically under infrared coarse graining. Physically, overdamped collective modes dissipate into the electronic continuum, continuously erasing temporal coherence of pairing fluctuations and preventing strong accumulation of superconducting correlations.
III.2 B. Infrared reorganization by slow collective modes
In the memory–dominated regime derived in Sec. II, collective dynamics is governed not by a single overdamped soft mode but by a continuum of slow relaxation modes characterized by the time–scale density of states $\rho(\lambda)$. After integrating out the relaxation reservoir, the retarded collective susceptibility takes the exact spectral form
$\chi_{R}^{-1}(\omega,\mathbf{q})=r+cq^{2}+\Sigma(\omega),\qquad\Sigma(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda-i\omega}.$ (29)
For a flat TDOS in the infrared, $\rho(\lambda\to 0)=\rho_{0}$, the effective polarization kernel in the Cooper channel inherits the infrared scaling
$\Pi(\omega)\propto\int_{0}^{\Lambda}d\lambda\,\rho_{0}\frac{1}{\lambda^{2}+\omega^{2}}\sim\frac{1}{|\omega|},\qquad|\omega|\to 0.$ (30)
Unlike the conventional Ohmic case, this singular behavior does not originate from Landau damping by fast electronic continua. Instead, it emerges from the accumulation of an extensive continuum of near–marginal relaxation modes encoded in a flat TDOS. The resulting kernel corresponds in the time domain to a universal long–memory response $K(t)\sim 1/t$.
Although both Eliashberg theory and the present memory–dominated regime generate infrared kernels scaling as $1/|\omega|$, their physical origins are fundamentally distinct. In Eliashberg theory the singularity reflects overdamped dissipation arising from coupling to a fast electronic bath [25, 26]. Here, by contrast, the same infrared scaling results from spectral accumulation of slow collective modes, producing intrinsically non–Markovian dynamics.
This dynamical reorganization qualitatively alters the Cooper channel: instead of marginal logarithmic enhancement characteristic of overdamped criticality, pairing fluctuations accumulate coherently across widely separated time scales. The slow–mode reservoir therefore converts the Cooper channel into a relevant infrared interaction.
III.3 C. Relevant infrared amplification of the Cooper channel
The infrared shell correction in the memory–dominated regime becomes
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}\ln(\Lambda/\omega^{\prime})}.$ (31)
Introducing $u=\ln(\Lambda/\omega^{\prime})$ yields
$dg\simeq\alpha g\,dl,$ (32)
with $\alpha>0$ proportional to the infrared TDOS weight.
The resulting RG flow reads
$\frac{dg}{dl}=\alpha g,$ (33)
with solution $g(l)=g_{0}e^{\alpha l}$.
The Cooper–channel interaction therefore grows algebraically under infrared coarse graining, signaling a relevant superconducting instability rather than a marginal one. This qualitative contrast with Eliashberg theory is illustrated in Fig. 2. A detailed shell RG derivation demonstrating the crossover from logarithmic to power-law pairing growth is presented in Appendix D.
Physically, the slow–mode reservoir retains dynamical correlations across widely separated time scales, allowing collective fluctuations to act repeatedly on electronic degrees of freedom and amplify intrinsic electronic pairing tendencies.
Superconductivity in strongly correlated systems thus emerges from infrared spectral organization of collective relaxation dynamics, rather than from finely tuned exchange of specific bosonic mediators.
@@
favoring local singlet formation.
More generally, correlated electronic systems possess an intrinsic short–range pairing scale $E_{\rm pair}$ generated purely by electronic interactions, independent of long–wavelength collective mediators. This intrinsic scale provides the microscopic seed for superconductivity.
@@
with $g_{\rm pair}\sim E_{\rm pair}^{-1}$. Critical dynamics does not create attraction; it amplifies a preexisting electronic pairing tendency through infrared spectral reorganization.
@@
Because single–particle propagators admit a relaxational spectral representation governed by the slow–mode TDOS, the Cooper bubble reduces to a continuum integral over relaxation rates,
@@
This nonanalytic enhancement reflects the accumulation of collective memory across widely separated time scales. Unlike conventional Eliashberg kernels, which are marginal in the infrared, the present Cooper channel becomes relevant due to the flat TDOS. The slow–mode reservoir therefore reorganizes the pairing problem into a memory–dominated critical regime.
@@
which yields
@@
up to numerical factors of order unity. The finite-temperature regularization of the infrared divergence, leading to $\Pi(T)\sim 1/T$, is derived explicitly in Appendix E.
@@
In the memory–dominated regime, long–time coherence is controlled by the infrared TDOS weight $\rho_{0}$. Because the pairing amplitude itself scales as $\Delta\sim E_{\rm pair}\,\rho_{0}$ [Eq. (39)], the density of phase-coherent pairs is likewise proportional to the same infrared spectral weight. Integrating out amplitude fluctuations in the presence of the slow–mode reservoir therefore yields
@@
The TDOS controls: (i) the infrared extent of the slow–mode reservoir, (ii) the dynamical amplification of intrinsic pairing tendencies, (iii) the superfluid stiffness and phase coherence, and (iv) the emergence of non-Markovian long–time response. Quantum criticality thus acts not merely as a source of fluctuations but as a dynamical amplifier governed by infrared spectral reorganization that simultaneously governs thermodynamic and dynamical phenomena in strongly correlated superconductors.
This unified picture yields a concrete strategy for experimental tests: systems that exhibit enhanced low–frequency TDOS weight (or equivalently, strong long–time correlations) are expected to display increased pairing susceptibility, elevated $T_{c}$, and reduced phase stiffness consistent with Uemura scaling. The resulting real-frequency spectral function generated by the memory-dominated self-energy, including its universal scaling form and infrared structure, is derived in Appendix F.
@@
The central result of this work is that infrared collective dynamics in strongly correlated quantum matter is organized by the structure of its relaxation-rate spectrum rather than solely by softening of a single order parameter. The time–scale density of states provides a quantitative measure of this organization. When the TDOS remains finite as $\lambda\to 0$, the system enters a memory–dominated regime characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic infrared response.
@@
This viewpoint clarifies the limits of conventional coarse-grained field theories. Away from criticality, the relaxation spectrum is gapped in the infrared, leading to finite memory times and justifying reduction to local-in-time damping and a small number of collective fields. Near correlated critical regimes, however, relaxation rates condense into a continuous spectrum extending toward zero. In this regime, internal modes cannot be integrated out: the full decay spectrum remains dynamically active across time scales. The breakdown of Markovian coarse-graining is therefore not a model-dependent artifact but a universal consequence of dynamical criticality.
@@
This observation highlights the structural difference from Hertz–Millis–type approaches. Conventional quantum critical theories assume overdamped local dynamics of a small set of bosonic modes, with dissipation controlled by Landau damping. In contrast, the present framework allows for spectral condensation of near-marginal relaxation channels. Criticality is reinterpreted not as single-mode softening but as reorganization of the entire decay spectrum. The TDOS thus replaces phenomenological damping coefficients with a microscopic spectral quantity governing infrared universality.
@@
A direct physical consequence of this spectral reorganization is qualitative enhancement of superconducting instabilities. Superconductivity need not rely on finely tuned bosonic mediators. Instead, intrinsic short-range electronic pairing tendencies are dynamically amplified by the slow-mode reservoir. The exponential suppression characteristic of weak-coupling BCS theory is replaced by algebraic scaling controlled by infrared spectral weight. In this sense, superconducting order corresponds to condensation within an extended slow-mode manifold rather than binding via a sharp bosonic excitation.
@@
A single infrared parameter, $\rho_{0}$, governs both thermodynamic and dynamical phenomena. The superconducting dome reflects truncation of long-time memory away from the optimal spectral reorganization point; Uemura scaling follows from the shared spectral origin of pairing and phase coherence; anomalous normal-state dynamics arises from the same non-Markovian kernel. Quantum criticality therefore acts as a dynamical amplifier linking transport, coherence, and pairing within a unified spectral framework.
@@
The relaxation-rate organization emphasized here is not merely a phenomenological assumption. Both Liouvillian evolution in open quantum systems and generalized memory-kernel dynamics naturally decompose into continua of decay modes characterized by complex relaxation eigenvalues. Whenever long-time correlations are present, Laplace-domain representations generically produce continuous relaxation spectra. A dense TDOS therefore reflects a universal mathematical property of dynamical systems with extended temporal correlations, rather than a model-specific construction.

### v3

# Memory-Dominated Quantum Criticality as a Universal Route to High-Temperature Superconductivity

@@ Abstract
A key limitation of many conventional approaches lies not in the use of collective fields themselves, but in the assumption that the infrared dynamics can be adequately described by a small number of soft bosonic modes with local-in-time damping [10, 11]. Such descriptions effectively close the theory within a finite set of order-parameter degrees of freedom. While this approximation is often adequate in weakly interacting systems, its validity becomes less clear in strongly correlated electronic matter, where experimentally relevant observables correspond to coarse-grained collective sectors that continuously exchange coherence and energy with a large number of internal degrees of freedom. In such situations the effective dynamics need not be governed by a finite set of collective modes, but may instead involve a broad spectrum of relaxation processes spanning many time scales.
Although the full many-body system evolves unitarily under a Hamiltonian, the effective dynamics of these collective sectors is more naturally described as that of an open subsystem. In this situation the appropriate dynamical generator is not simply the Hamiltonian itself but the effective Liouvillian governing relaxation and memory processes of the reduced density matrix [12, 13, 14]. The long-time behavior of the system is therefore controlled by the spectrum of relaxation eigenvalues of this Liouvillian operator. From this perspective, dynamical criticality need not correspond to the softening of a single excitation mode but may instead arise from a reorganization of the entire relaxation-rate spectrum.
Rather than reducing the dynamics to a few collective coordinates, one must therefore consider the possibility that the Liouvillian spectrum develops an extended continuum of near-marginal decay modes. Such spectral accumulation naturally generates a broad distribution of relaxation time scales and invalidates the assumption of Markovian coarse-graining near criticality. The relevant organizing structure of the infrared theory is then not a small set of overdamped bosons, but the time–scale density of states (TDOS) of collective relaxation modes.
@@
We demonstrate that the infrared scaling of the relaxation-rate density of states, $\rho(\lambda)\sim\lambda^{\alpha}$, naturally classifies universality classes of dynamical critical behavior. Within the relaxation-spectrum framework, this scaling directly determines the low-frequency structure of the retarded susceptibility. While the conventional Hertz–Millis kernel $\mathrm{Im}\,\chi^{-1}\propto|\omega|$ corresponds to the case $\alpha=1$, arising microscopically from particle–hole polarization of a Fermi liquid, a finite TDOS at the origin ($\alpha=0$) produces a marginal nonanalytic kernel proportional to $\mathrm{sgn}(\omega)$. In the time domain this corresponds to a universal long–memory response $K(t)\sim 1/t$. The latter regime reflects a fundamentally different physical mechanism: rather than overdamped order-parameter fluctuations generated by a fermionic bath, the infrared dynamics is governed by the collective organization of the relaxation spectrum itself. This memory-dominated regime therefore defines a distinct universality class controlled by the structure of the TDOS.
@@
the collective susceptibility follows directly from the spectral representation in Eq. (16).
The dissipative part of the kernel is
@@
Because the response function is causal, the reactive component is not independent but is related to the dissipative part through the Kramers–Kronig relation. The real part therefore inherits the same infrared scaling structure and, depending on the TDOS exponent, produces logarithmic or subleading corrections to the static stiffness.
It is convenient to express the long–wavelength response in the scaling form
@@
where the TDOS scaling directly implies the identification $s=\alpha$.
While this form resembles the conventional Hertz–Millis representation, its physical interpretation is different. In the present framework the exponent $s$ does not originate from fermionic damping of a single order–parameter mode but instead reflects the infrared organization of the entire relaxation spectrum.
The appropriate renormalization viewpoint is therefore naturally formulated in terms of temporal coarse–graining. Integrating out fast relaxation processes with decay rates $\lambda>\Lambda/b$ corresponds to eliminating short time scales $t\lesssim b^{-1}\Lambda^{-1}$. Under this temporal renormalization step the remaining effective theory is characterized by a renormalized TDOS
$\rho_{b}(\lambda)=b^{\alpha}\rho(b\lambda),$ (20)
so that the infrared scaling exponent $\alpha$ defines the universality class of the relaxation spectrum. Infrared universality classes are therefore classified directly by the low–energy structure of the relaxation–rate density.
A particularly important case occurs when the TDOS remains finite at the origin,
$\rho(\lambda)\xrightarrow{\lambda\to 0}\rho_{0}.$ (21)
In this flat–TDOS regime the dissipative kernel simplifies to
@@
while the reactive component develops a logarithmic infrared singularity,
@@
Collective infrared dynamics is therefore governed by the marginal non–Markovian kernel
@@
which corresponds in the time domain to a universal long–memory response $K(t)\sim 1/t$, derived explicitly in Appendix C.
This flat–TDOS fixed point represents a dynamical regime in which the infrared behavior is governed by the accumulation of a continuum of slow decay modes. Rather than arising from overdamped fluctuations of a single soft order parameter, collective relaxation is controlled by the spectral organization of the full relaxation spectrum.
@@
The dynamical framework developed in the previous section implies that strongly correlated quantum matter can self–organize into a regime dominated by a continuum of slow collective modes. Instead of a single overdamped bosonic mediator, the infrared dynamics is governed by the full relaxation spectrum characterized by the time–scale density of states $\rho(\lambda)$.
This reorganization of the collective spectrum qualitatively alters the structure of the Cooper channel. In conventional approaches superconductivity is typically attributed to the exchange of a specific bosonic “pairing glue”. In contrast, in the present framework superconductivity emerges from the dynamical amplification of pairing correlations generated by the slow–mode reservoir itself.
In this section we derive the infrared structure of the Cooper channel in this regime and contrast it with the conventional Eliashberg mechanism.
III.1 A. Marginal pairing from overdamped collective dynamics
Near metallic quantum critical points, electrons often couple to overdamped collective modes whose retarded propagator takes the form
@@
Here $D_{R}$ denotes the effective propagator of a coarse-grained order-parameter (particle–hole) collective mode in the Hertz–Millis sense: the $|\omega|$ term is generated by the electron–hole polarization of the metallic continuum (Landau damping) [3, 4]. This object should be distinguished from the Cooper-channel polarization $\Pi(\omega)$ introduced below, which diagnoses the pairing instability itself.
@@
The linear frequency dependence reflects Landau damping due to decay of collective excitations into the electronic continuum. After momentum integration, the pairing kernel entering the Cooper channel behaves approximately as
$D_{R}(\omega)\sim\frac{1}{|\omega|}.$ (26)
The linearized gap equation may be written schematically as
$\Delta(\omega)=\int_{T}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}\,g(\omega-\omega^{\prime})\,\Delta(\omega^{\prime}),$ (27)
where $\Lambda$ is the ultraviolet cutoff and $T$ provides the infrared cutoff.
To analyze the infrared flow we eliminate fast frequency modes with $\omega^{\prime}\in[\Lambda/b,\Lambda]$, corresponding to coarse–graining of short time scales. The ladder diagram gives the leading correction
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}.$ (28)
The shell integral yields
$\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}=\ln b.$ (29)
Introducing the RG scale $l=\ln b$ gives
$\frac{dg}{dl}=Cg^{2}.$ (30)
@@
The Cooper interaction therefore grows only logarithmically under infrared coarse graining. In RG language the pairing interaction is marginal. Physically, overdamped collective modes dissipate into the electronic continuum and repeatedly lose phase coherence. As a result superconducting correlations accumulate only logarithmically as the energy scale is reduced.
III.2 B. Collective relaxation spectrum and TDOS representation
Strongly correlated quantum matter, however, can develop a much richer infrared structure. Instead of a single overdamped bosonic excitation, the system may host a dense continuum of slow collective modes.
The collective dynamics can then be represented by the TDOS $\rho(\lambda)$, which measures the density of relaxation rates $\lambda$. Projecting the dynamics onto the slow collective sector yields the spectral representation
$\Sigma(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda-i\omega}.$ (31)
The polarization kernel entering the Cooper channel becomes
$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}.$ (32)
Here $\Pi(\omega)$ represents the polarization of the Cooper channel generated by the collective relaxation spectrum. It should be distinguished from the conventional bosonic polarization that appears in propagator renormalization. In the present framework $\Pi(\omega)$ directly measures the dynamical pair susceptibility produced by the slow collective modes.
For a flat TDOS in the infrared,
$\rho(\lambda\to 0)=\rho_{0},$ (33)
the polarization acquires a universal singular form
$\Pi(\omega)\sim\frac{\rho_{0}}{|\omega|}.$ (34)
In the time domain this corresponds to a long–memory kernel
$K(t)\sim\frac{1}{t}.$ (35)
Importantly, this singular behavior does not originate from Landau damping into a fast electronic bath. Instead it arises from the spectral accumulation of slow relaxation modes themselves.
III.3 C. Infrared amplification of pairing correlations
The Cooper susceptibility takes the form
$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{pair}}-\Pi(\omega).$ (36)
Because $\Pi(\omega)$ diverges as $|\omega|\to 0$, pairing correlations are strongly enhanced as the energy scale is lowered.
To understand the infrared flow more generally, consider a TDOS scaling
$\rho(\lambda)\sim\lambda^{\eta}.$ (37)
Substituting this form into the polarization integral yields
$\Pi(\omega)\sim\int d\lambda\frac{\lambda^{\eta}}{\lambda^{2}+\omega^{2}}\sim|\omega|^{-(1-\eta)}.$ (38)
The effective pairing kernel therefore scales as
$g_{\mathrm{eff}}(\omega)\sim|\omega|^{-(1-\eta)}.$ (39)
Under RG rescaling $\omega\rightarrow\omega/b$ the kernel transforms as
$g^{\prime}=b^{1-\eta}g.$ (40)
Introducing $l=\ln b$ gives the tree–level RG flow
$\frac{dg}{dl}=(1-\eta)g.$ (41)
For $\eta<1$ the Cooper interaction becomes relevant in the infrared. Including the usual ladder correction yields
$\frac{dg}{dl}=(1-\eta)g+Cg^{2}.$ (42)
The linear term dominates the early RG flow, leading to algebraic growth
$g(l)=g_{0}e^{(1-\eta)l}.$ (43)
Physical interpretation. The qualitative difference between the Eliashberg regime and the present memory–dominated regime originates from the structure of the collective spectrum, as illustrated schematically in Fig. 2.
In the Eliashberg case, each logarithmic shell contributes independently, leading to marginal enhancement of pairing correlations. In contrast, a flat TDOS corresponds to a dense continuum of near–marginal relaxation modes. Fluctuations across widely separated time scales remain dynamically correlated, allowing pairing correlations to accumulate coherently.
Superconductivity therefore emerges as a dynamical instability of the collective relaxation spectrum itself, rather than from the exchange of a specific bosonic mediator.
@@
which favors local singlet formation. This interaction provides a natural microscopic origin of electronic pairing in systems near Mott localization.
More generally, interacting electrons are described by a generic four–fermion interaction
$H_{\rm int}=\sum_{k,k^{\prime},q}V(q)\,c^{\dagger}_{k+q\sigma}c^{\dagger}_{k^{\prime}-q\sigma^{\prime}}c_{k^{\prime}\sigma^{\prime}}c_{k\sigma}.$ (46)
This interaction contains multiple collective channels. In conventional diagrammatic approaches these channels are analyzed through resummations such as particle–hole bubble diagrams or particle–particle ladder diagrams. Rather than focusing on a specific mediator, the present framework emphasizes the dynamical organization of the collective spectrum itself.
While the four–fermion interaction generically allows multiple ordering channels, their ordering tendencies are not equivalent. Particle–hole instabilities such as charge or spin density waves typically require specific momentum structure, most notably Fermi surface nesting or commensurate wave vectors determined by the lattice. Such kinematic constraints restrict the available phase space for density–wave ordering.
Superconductivity, by contrast, corresponds to condensation in the particle–particle channel at zero total momentum,
$\Delta=\langle c_{k\uparrow}c_{-k\downarrow}\rangle,$ (47)
and therefore does not rely on special nesting conditions of the Fermi surface. Consequently the Cooper channel can efficiently accumulate infrared spectral weight generated by collective fluctuations.
@@
where $g_{\rm pair}$ characterizes the intrinsic electronic pairing scale $E_{\rm pair}$ generated by short–range correlations. Diagrammatically $\Pi(\omega)$ corresponds to the Cooper bubble, which resums particle–particle ladder diagrams.
In weakly interacting systems the pairing instability is exponentially suppressed as in conventional BCS theory. Near correlated criticality, however, the situation changes qualitatively. Because the relaxation spectrum becomes dense near $\lambda\to 0$, an extensive reservoir of slow collective modes emerges. This slow–mode reservoir reorganizes the infrared dynamics of the electronic system and strongly enhances the Cooper channel.
Within the present framework critical dynamics does not create attraction. Instead it amplifies a preexisting electronic pairing tendency through infrared spectral reorganization. The intrinsic pairing scale $E_{\rm pair}$ therefore acts as the microscopic seed, while the slow collective spectrum dynamically amplifies the Cooper instability.
@@
Because single–particle propagation inherits the relaxational spectral structure generated by the slow–mode reservoir, the Cooper bubble reduces to a continuum integral over relaxation rates,
@@
This nonanalytic enhancement reflects the accumulation of collective memory across widely separated time scales. Physically, the flat TDOS implies that an extensive set of slow collective modes contributes coherently to the Cooper channel. Unlike conventional Eliashberg kernels, which remain marginal in the infrared, the present Cooper kernel becomes strongly enhanced by the infrared spectral weight of the slow–mode reservoir. The pairing interaction is therefore dynamically amplified by the collective memory of the system.
@@
which yields the algebraic transition scale
@@
up to numerical factors of order unity. The finite-temperature regularization of the infrared divergence, leading to $\Pi(T)\sim 1/T$, is derived explicitly in Appendix D.
@@
In the memory–dominated regime, long–time coherence is controlled by the infrared TDOS weight $\rho_{0}$. Because the pairing amplitude itself scales as $\Delta\sim E_{\rm pair}\,\rho_{0}$, the density of phase-coherent pairs is likewise proportional to the same infrared spectral weight. Integrating out amplitude fluctuations in the presence of the slow–mode reservoir therefore yields
@@
The density of relaxation rates controls: (i) the infrared extent of the slow–mode reservoir, (ii) the dynamical amplification of intrinsic pairing tendencies, (iii) the superfluid stiffness and phase coherence, and (iv) the emergence of non-Markovian long–time response. Quantum criticality thus acts not merely as a source of fluctuations but as a dynamical amplifier governed by infrared spectral reorganization that simultaneously governs thermodynamic and dynamical phenomena in strongly correlated superconductors.
This unified picture yields a concrete strategy for experimental tests: systems that exhibit enhanced low–frequency TDOS weight (or equivalently, strong long–time correlations) are expected to display increased pairing susceptibility, elevated $T_{c}$, and reduced phase stiffness consistent with Uemura scaling. The resulting real-frequency spectral function generated by the memory-dominated self-energy, including its universal scaling form and infrared structure, is derived in Appendix E.
Figure 4: Universality map relating the relaxation-rate spectrum to infrared collective dynamics and superconducting pairing. The TDOS exponent $\alpha$ determines the scaling of the polarization kernel $\Pi(\omega)\sim|\omega|^{\alpha-1}$. The Ohmic Hertz–Millis theory corresponds to the marginal point $\alpha=1$, while a flat TDOS ($\alpha=0$) produces memory-dominated critical dynamics with $\Pi(\omega)\sim 1/|\omega|$ and enhanced pairing.
@@
The central result of this work is that infrared collective dynamics in strongly correlated quantum matter is organized by the structure of its relaxation-rate spectrum rather than solely by the softening of a single order parameter. The time–scale density of states provides a quantitative measure of this organization. When the TDOS remains finite as $\lambda\to 0$, the system enters a memory–dominated regime characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic infrared response functions.
@@
This viewpoint clarifies the limits of conventional coarse-grained field theories. Away from criticality, the relaxation spectrum is gapped in the infrared, leading to finite memory times and justifying reduction to local-in-time damping and a small number of collective fields. Near correlated critical regimes, however, relaxation rates condense into a continuous spectrum extending toward zero. In this regime internal modes cannot be integrated out: the full decay spectrum remains dynamically active across time scales. The breakdown of Markovian coarse-graining is therefore not a model-dependent artifact but a universal consequence of dynamical criticality.
@@
This observation highlights a structural difference from Hertz–Millis–type approaches. Conventional quantum critical theories assume overdamped local dynamics of a small set of bosonic modes, with dissipation controlled by Landau damping. In contrast, the present framework allows for spectral condensation of near-marginal relaxation channels. Criticality is reinterpreted not as single-mode softening but as a reorganization of the entire decay spectrum. The TDOS therefore replaces phenomenological damping coefficients with a microscopic spectral quantity governing infrared universality.
This structure is summarized schematically in Fig. 4. The figure illustrates how the infrared universality class of the collective dynamics is determined by the exponent of the relaxation-rate density of states. For a TDOS scaling $\rho(\lambda)\sim\lambda^{\alpha}$, the polarization kernel assumes the universal infrared form
$\Pi(\omega)\sim|\omega|^{\alpha-1}.$ (65)
Different dynamical regimes therefore emerge as special cases of the same spectral organization. For $\alpha>1$ the density of slow modes vanishes in the infrared, so that $\Pi(\omega)\rightarrow 0$ and the collective dynamics reduces to weakly damped propagating modes with $\chi^{-1}\simeq r+cq^{2}$. The marginal case $\alpha=1$ corresponds to Ohmic dissipation, reproducing the Hertz–Millis form $\chi^{-1}\simeq r+cq^{2}+\gamma|\omega|$. Most importantly, when the TDOS approaches a finite constant ($\alpha=0$), a dense continuum of near-marginal relaxation modes emerges. In this regime the polarization acquires the singular form $\Pi(\omega)\sim 1/|\omega|$, producing long-time memory kernels $K(t)\sim 1/t$ and the non-Markovian infrared dynamics discussed throughout this work.
The lower panel of Fig. 4 illustrates the direct consequence of this spectral organization for superconductivity. Within the Thouless criterion $\chi_{\Delta}^{-1}(0,T)=g_{\rm pair}^{-1}-\Pi(T)$, Ohmic dynamics leads to the familiar logarithmic kernel $\Pi(T)\sim\ln(\Omega/T)$ and the exponentially small BCS transition temperature. In contrast, memory-dominated dynamics with a flat TDOS generates the stronger scaling $\Pi(T)\sim C/T$, converting moderate microscopic pairing into an algebraically enhanced instability $T_{c}\sim\rho_{0}E_{\rm pair}$. The figure therefore provides a compact universality map linking the relaxation-rate spectrum, infrared collective dynamics, and the resulting pairing scale.
@@
A direct physical consequence of this spectral reorganization is the enhancement of superconducting instabilities. Superconductivity need not rely on finely tuned bosonic mediators. Instead, intrinsic short-range electronic pairing tendencies are dynamically amplified by the slow-mode reservoir. The exponential suppression characteristic of weak-coupling BCS theory is replaced by algebraic scaling controlled by infrared spectral weight. In this sense, superconducting order corresponds to condensation within an extended slow-mode manifold rather than binding via a sharp bosonic excitation.
@@
A single infrared parameter, $\rho_{0}$, governs both thermodynamic and dynamical phenomena. The superconducting dome reflects truncation of long-time memory away from the optimal spectral reorganization point; Uemura scaling follows from the shared spectral origin of pairing and phase coherence; and anomalous normal-state dynamics arises from the same non-Markovian kernel. Quantum criticality therefore acts as a dynamical amplifier linking transport, coherence, and pairing within a unified spectral framework.
Liouvillian description of infrared collective dynamics.
@@
The relaxation-rate organization emphasized here is not merely a phenomenological assumption. Both Liouvillian evolution in open quantum systems and generalized memory-kernel dynamics naturally decompose into continua of decay modes characterized by complex relaxation eigenvalues. Whenever long-time correlations are present, Laplace-domain representations generically produce continuous relaxation spectra. A dense TDOS therefore reflects a universal mathematical property of dynamical systems with extended temporal correlations rather than a model-specific constructione, is derived in Appendix F.
This viewpoint is closely related to the distinction between closed and effective open descriptions of many-body dynamics. While the complete microscopic system evolves unitarily under a Hamiltonian, the experimentally accessible collective sector behaves as a coarse-grained subsystem continuously exchanging coherence and information with the remaining degrees of freedom. From the perspective of such a subsystem, the relevant generator of dynamics is not the Hamiltonian itself but the effective Liouvillian governing relaxation and memory processes. The infrared behavior of strongly correlated matter is therefore naturally organized by the spectrum of relaxation eigenvalues rather than solely by the spectrum of excitation energies. In this sense the TDOS plays a role analogous to a Liouvillian spectral density, providing a dynamical counterpart to the Lehmann representation familiar from Hamiltonian many-body theory.